# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ Abstract
Multivariate Hawkes process provides a powerful framework for modeling temporal dependencies and event-driven interactions in complex systems. While existing methods primarily focus on uncovering causal structures among observed subprocesses, real-world systems are often only partially observed, with latent subprocesses posing significant challenges. In this paper, we show that continuous-time event sequences can be represented by a discrete-time model as the time interval shrinks, and we leverage this insight to establish necessary and sufficient conditions for identifying latent subprocesses and the causal influences. Accordingly, we propose a two-phase iterative algorithm that alternates between inferring causal relationships among discovered subprocesses and uncovering new latent subprocesses, guided by path-based conditions that guarantee identifiability. Experiments on both synthetic and real-world datasets show that our method effectively recovers causal structures despite the presence of latent subprocesses.
@@ 1 Introduction
Understanding causality in complex systems is essential across diverse scientific and practical domains, including social networks (zhou2013learning, ), neuroscience (bonnet2022neuronal, ), and finance (hawkes2018hawkes, ). Multivariate Hawkes processes (hawkes1971spectra, ; laub2015hawkes, ), with their ability to model temporal dependencies and event-driven interactions, have emerged as powerful tools for capturing these dynamics. A majority of existing approaches to learning Hawkes processes (xu2016learning, ; eichler2017graphical, ; salehi2019learning, ; ide2021cardinality, ) are grounded in Granger causality (kim2011granger, ) and rely on MM-based maximum likelihood estimation (veen2008estimation, ) for model fitting. To mitigate the reliance on high-resolution event timestamps, several methods adopt coarse-grained representations by discretizing the timeline into bins (Kirchner_2016, ; shlomovich2022parameter, ; qiao2023structural, ; koyama2025coarse, ) and maximizing the likelihood of discretized processes.
However, all of the aforementioned methods operate under the sufficiency assumption—that all task-relevant subprocesses are fully observed—and focus on uncovering causal structures exclusively among these observed subprocesses. In practice, this assumption often fails, as many real-world systems are only partially observable, with some components remaining entirely unmeasured. For instance, in point process analysis within neuroscience, technical limitations in neural recording result in the temporal spike activity of many neurons going unrecorded, despite directly influencing the activity of observed neurons (huang2015effects, ), thereby obscuring the underlying causal structure. While prior work addressed missing event data from observed subprocesses by sampling over their posterior distributions (shelton2018hawkes, ), such approach still fails to handle truly latent subprocesses—those for which no events are observed at all. The presence of such latent subprocesses poses significant challenges to reliable causal discovery, particularly when they serve as latent confounders of other subprocesses.
In this paper, we tackle the largely unexplored problem of causal discovery in Hawkes processes under partial observability, focusing on identifying causal structures among subprocesses even in the presence of latent subprocesses. Leveraging a discrete-time representation, we propose a principled framework based on rank constraints of cross-covariance matrices, enabling the identification of latent subprocesses and causal influences. Our main contributions are:
  • We provide, to the best of our knowledge, the first framework for identifying latent subprocesses in continuous-time event sequences, addressing an important but underexplored problem in both Hawkes processes and discrete time series analysis.
@@
  • We establish that multivariate Hawkes processes can be represented by linear causal models over discretized variables, and derive necessary and sufficient conditions for identifying latent subprocesses and inferring causal influences.
@@
  • We develop a two-phase iterative algorithm that alternates between causal structure recovery and latent subprocess discovery, without requiring prior knowledge of latent components. It stops inference when identifiability conditions are no longer satisfied, ensuring robustness in practice.
@@ 2.1 Multivariate Hawkes Process
A multivariate Hawkes process is a class of self-exciting point processes that model temporal dependencies among events. It consists of a set of counting processes $\mathcal{N}_{\mathcal{G}}\coloneq\{N_{i}\}^{l}_{i=1}$, where $N_{i}(t)$ denotes the number of type-$i$ events that have occurred up to time $t$. The formal definition is provided below, and a detailed exposition of its derivation and key properties is available in Appendix˜B. An illustrative example of such a process is shown in Fig. 1(a).
Definition 2.1 (Multivariate Hawkes Process (hawkes1971spectra, ; laub2015hawkes, )).
A multivariate Hawkes process is a collection of $l$ counting processes such that for each $i\in\{1,\dots,l\}$, the intensity of counting process $N_{i}$ that governs the event-triggering behavior can be written as:
@@
where $\mu_{i}$ is the background intensity and $\sum_{j=1}^{l}\int_{0}^{t}\phi_{ij}(t-s)dN_{j}(s)$ is the endogenous intensity measuring the peer influence. Piecewise-continuous function $\phi_{ij}(s)\geq 0,\forall s\in(0,\infty)$ is called excitation function, which measures decay in the influence of historical type-$j$ events on the subsequent type-$i$ events. Typically, we consider (strictly) stationarity to avoid explosion, which means that for the influence matrix $\Phi\in\mathbb{R}^{l\times l}$ with entries $\Phi_{ij}=\int_{0}^{\infty}\phi_{ij}(s)ds$, the spectral radius $\rho(\Phi)<1$ (jovanovic2015cumulants, ; bacry2016first, ).
We are interested in identifying, for each subprocess (i.e., counting process) $N_{i}$, the minimal set of subprocesses $\mathcal{P}_{\mathcal{G}}\subseteq\mathcal{N}_{\mathcal{G}}$ such that the intensity function $\lambda_{i}(t)$ depends only on the historical events of the subprocesses in $\mathcal{P}_{\mathcal{G}}$, and not on those of the remaining subprocesses. Formally, this corresponds to $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$ for each $N_{j}\in\mathcal{P}_{\mathcal{G}}$, and zero otherwise. In this case, $N_{i}$ is said to be locally independent (didelez2008graphical, ) of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$, precisely because the integral is zero for every $N_{j}\in\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$.
@@ 2.2 Model Definition
To formalize our framework, we define a graphical causal model for multivariate Hawkes processes, where nodes represent subprocesses and directed edges correspond to nonzero excitation functions. We then provide a formal definition of causal effects for subsequent analysis. The overall goal is to identify latent subprocesses and infer causal relations among both observed and latent subprocesses.
@@
Let $\mathcal{G}\coloneq(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}})$ be a directed graph, where each node $N_{i}\in\mathcal{N}_{\mathcal{G}}$ represents a subprocess in a multivariate Hawkes process. A directed edge $E_{ij}\in\mathcal{E}_{\mathcal{G}}$ exists if and only if the excitation function satisfies $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$. The node set $\mathcal{N}_{\mathcal{G}}$ consists of $p+q$ nodes, comprising $p$ observed nodes $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$ and $q$ latent nodes $\mathcal{L}_{\mathcal{G}}\coloneqq\{L_{i}\}_{i=1}^{q}$, which correspond to the observed and latent subprocesses, respectively, in the underlying multivariate Hawkes process.
It is important to note that the PO-MHP model defined in Definition˜2.2 is not restricted to directed acyclic graphs. In fact, directed cycles and self-loops may naturally arise, which are typically challenging to analyze (claassen2023establishing, ). Interestingly, two subprocesses $N_{i}$ and $N_{j}$ may form a directed cycle if there exist directed paths from $N_{i}$ to $N_{j}$ and from $N_{j}$ back to $N_{i}$. Any subprocess $N_{i}$ has self loop if $\int_{0}^{t}\phi_{ii}(t-s)\,dN_{i}(s)>0$. Furthermore, we allow for the presence of latent subprocesses, and directed edges may exist between any pair of subprocesses, whether observed or latent. In this setting, both $N_{i}$ and $N_{j}$ may be either observed or latent. To the best of our knowledge, it is the first work to investigate such a general structure in Hawkes processes.
Definition 2.3 (Causal Effect).
For any two subprocesses $N_{i}$ and $N_{j}$ (observed or latent) in a multivariate Hawkes process, if there exists a directed path from $N_{i}$ to $N_{j}$, then $N_{j}$ is said to be a effect of $N_{i}$, and $N_{i}$ is said to be a cause of $N_{j}$.
Definition 2.4 (Parent Cause Set).
@@
Proposition 2.5 (Parent Cause Set and Local Independence).
A subprocess $N_{i}$ is locally independent (as defined in Section˜2.1) of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$ if and only if $\mathcal{P}_{\mathcal{G}}$ is the parent cause set of $N_{i}$ in $\mathcal{N}_{\mathcal{G}}$.
Figure 1: Illustration of a multivariate Hawkes process with three subprocesses $N_{1},N_{2},N_{3}$. (a) A point process representation, where the continuous timeline is partitioned into intervals of length $\Delta$. (b) The corresponding summary causal graph, which is the central object of this paper. Each node represents a subprocess, with causal relations $N_{1}\leftarrow N_{2}\rightleftharpoons N_{3}$, and self-loops on all nodes. (c) The window causal graph, depicting the underlying time-lagged causal mechanism. Each node denotes the count in a time interval $\Delta$, modeled as a weighted sum of lagged parent nodes and an uncorrelated noise, as in Eq. 2. Note: This paper focuses on cases where some subprocesses are latent.
@@ 3.1 From Continuous-Time to Discrete-Time Representation
Directly inferring the causal structure among subprocesses in a multivariate Hawkes process modeled by Eq. 1, is challenging, particularly in the presence of latent subprocesses, as Eq. 1 characterizes a continuous-time stochastic process in which $\lambda_{i}(t)$ denotes the expected instantaneous rate of type-$i$ events at time $t$, conditioned on the event history. To address this, we discretize the timeline into bins (Kirchner_2016, ; shlomovich2022parameter, ; qiao2023structural, ; koyama2025coarse, ) and establish a connection between continuous-time Hawkes processes and discrete-time linear autoregressive models, which facilitates both the identification of latent subprocesses and the inference of the underlying causal structure.
Theorem 3.1 (Representation of the Multivariate Hawkes Process as a Linear Autoregressive Model).
Let $\mathcal{N}_{\mathcal{G}}\coloneq\{N_{i}\}^{l}_{i=1}$ be a stationary multivariate Hawkes process with background intensities $\{\mu_{i}\}_{i=1}^{l}$ and piecewise-continuous excitation functions $\{\phi_{ij}(s)\geq 0,\forall s\in(0,\infty)\}_{i,j=1}^{l}$. Define the discretized event count in the $n$-th time window of size $\Delta\in(0,\delta)$ as
@@
where $\delta$ is a constant depending on the moment structure of the process. Then, as $\Delta\rightarrow 0$, the Hawkes process admits the following linear autoregressive representation:
@@
where $\theta^{(0)}_{i}=\Delta\cdot\mu_{i}$ is the background parameter, $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)ds$ are the excitation coefficients, and $\varepsilon_{i}^{(n)}$ denotes the $n$-th realization of a serially uncorrelated white noise sequence.
Unlike the stochastic formulation in Eq. 1, Theorem˜3.1 reveals a deterministic relationship between the count in the current bin $N_{i}^{(n)}$ and counts in lagged bins $\{N_{j}^{(n-k)}\}_{j\in\{1,\dots,l\}}^{k\in\{1,\dots,n\}}$ (referred to as variables hereafter). Each current variable $N_{i}^{(n)}$ can be expressed as a weighted sum of all lagged variables, together with an uncorrelated noise term. This discrete-time representation enables the analysis of variable dependencies without modeling the entire continuous-time stochastic subprocesses, and allows for causal structure inference among subprocesses by leveraging the information encoded in the discretized form. As illustrated in Fig. 1, a directed edge $N_{2}\rightarrow N_{1}$ in the summary causal graph corresponds to the presence of directed edges from all lag variables $\{N_{2}^{(j)}\}_{j=0}^{n-1}$ to the current variable $N_{1}^{(n)}$, for all $n\in\mathbb{N}_{0}$. The proof of Theorem˜3.1 is provided in Appendix˜G.
In practice, it is unnecessary to consider all lagged variables for $k=1,\dots,n$ across the entire process history. This is because the excitation function $\phi_{ij}(s)$, which serves as a decay kernel, typically has finite support. For sufficiently large lags $k$, the corresponding excitation coefficient $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)\,ds$ approaches zero, implying that lagged variables $N_{j}^{(n-k)}$ beyond a certain point have negligible influence on $N_{i}^{(n)}$ (Kirchner_2016, ; kirchner2016hawkes, ). We refer to the smallest such cutoff point as the number of effective lag variables. In our analysis and practice, we consider at most $m$ lag variables, where $m$ is chosen to be greater than or equal to the number of effective lags. This approach is also standard in time series modeling, where a finite window facilitates sample collection and estimation (hyvarinen2010estimation, ; peters2013causal, ). However, prior works typically assume acyclic summary graphs and understate latent variables. Moreover, it is sufficient to choose a small bin width $\Delta$ relative to the typical support of the excitation function (Kirchner_2016, ; koyama2025coarse, ). For an empirical illustration of this point, refer to Section˜Q.3.

### v2

@@ Abstract
Multivariate Hawkes process provides a powerful framework for modeling temporal dependencies and event-driven interactions in complex systems. While existing methods primarily focus on uncovering causal structures among observed subprocesses, real-world systems are often only partially observed, with latent subprocesses posing significant challenges. In this paper, we show that continuous-time event sequences can be represented by a discrete-time causal model as the time interval shrinks, and we leverage this insight to establish necessary and sufficient conditions for identifying latent subprocesses and the causal influences. Accordingly, we propose a two-phase iterative algorithm that alternates between inferring causal relationships among discovered subprocesses and uncovering new latent subprocesses, guided by path-based conditions that guarantee identifiability. Experiments on both synthetic and real-world datasets show that our method effectively recovers causal structures despite the presence of latent subprocesses.
@@ 1 Introduction
Understanding causality in complex systems is essential across diverse scientific and practical domains, including social networks (Zhou et al., 2013), neuroscience (Bonnet et al., 2022), and finance (Hawkes, 2018). Multivariate Hawkes processes (Hawkes, 1971), with their ability to model temporal dependencies and event-driven interactions, have emerged as powerful tools for capturing these dynamics. A majority of existing approaches (Xu et al., 2016; Salehi et al., 2019; Idé et al., 2021) learns Hawkes processes by using Granger causality (Kim et al., 2011) and fitting continuous-time event sequences via maximum likelihood (Veen and Schoenberg, 2008). Another line of work reduces reliance on high-resolution event timestamps by performing likelihood-based estimation directly on pre-binned counts (Shlomovich et al., 2022; Cai et al., 2022; Qiao et al., 2023).
Almost all existing methods, including the two lines above, implicitly assume causal sufficiency: all task-relevant subprocesses (i.e., event sequences) are fully observed, and the goal is to uncover causal structure only among those observed subprocesses. In practice this assumption is often violated: many real-world systems are only partially observable, with some event sequences entirely unmeasured. For example, in neuroscience, limitations of neural recording leave many neurons unobserved even though they influence recorded neurons (Huang, 2015), obscuring the true causal structure. Shelton et al. (2018) proposes a method to impute missing event times within an observed subprocess via posterior sampling, but this does not handle entirely unobserved subprocesses unless one specifies their existence and number in advance—an impractical requirement. Identifying such latent subprocesses is crucial for reliable causal discovery, particularly when they act as confounders.
In this paper, we target the general scenario of partial observability: we seek to recover the causal structure among both observed and latent subprocesses without prior knowledge about whether latent subprocesses exist, how many there are, or where they connect. We first show that, as the time interval shrinks, the multivariate Hawkes process admits a discrete-time linear causal representation. Leveraging second-order (cross-covariance) statistics of this representation, we show that the causal graph is virtually identifiable, without prior knowledge of latent subprocesses, when each latent subprocess has suitable observed surrogates. Our main contributions are:
  • To the best of our knowledge, we provide the first principled framework that identifies latent subprocesses and recovers causal structure in continuous-time event sequences without prior knowledge of the existence or number of latent subprocesses.
@@
  • By showing that multivariate Hawkes processes can be represented by a linear causal model over discretized variables, we derive necessary and sufficient conditions for identifying latent subprocesses and inferring causal influences.
@@
  • We develop a two-phase iterative algorithm that alternates between causal-structure recovery and latent-subprocess discovery, using rank tests on cross-covariance matrices of observed discretized variables; it requires no prior knowledge of latent components.
2 Related Work
Related work relevant to this paper are two areas: Hawkes processes and causal discovery. We briefly review the most relevant works here; a more detailed discussion is deferred to Appendix˜A.
Hawkes Processes. Hawkes processes provide a flexible framework for modeling temporal dependencies among events (Hawkes, 1971; Laub et al., 2015). Existing methods for learning Hawkes structures from continuous-time data predominantly rely on likelihood-based estimation, using either parametric kernels (e.g., exponential, power-law) (Xu et al., 2016) or nonparametric procedures (Lewis and Mohler, 2011), often combined with sparsity regularization (Zhou et al., 2013; Idé et al., 2021). The NPHC method (Achab et al., 2018) estimates integrated kernels via moment-matching, providing a nonparametric alternative. These approaches are effective when all subprocesses are observed, but they do not account for latent components. When only binned event counts are available, another line of work fits Hawkes models by maximizing likelihood over bin counts (Shlomovich et al., 2022; Cai et al., 2022; Qiao et al., 2023), but such methods again presuppose full observability. In contrast, although our framework also operates on discretized data, it departs from likelihood-based fitting: by leveraging the autoregressive representation of Hawkes processes, we exploit low-rank patterns in cross-covariances, thus enabling the identification of latent subprocesses.
Causal Discovery. Existing causal discovery methods are primarily designed for settings where variables follow deterministic relations, rather than stochastic event-driven dynamics such as Hawkes processes. Classical approaches for i.i.d. data include constraint-based (Spirtes et al., 2001), score-based (Chickering, 2002), and functional approaches (Shimizu et al., 2006). Extensions to handle latent variables include rank-based methods that identify hidden structures in linear models (Huang et al., 2022; Dong et al., 2023). While these allow latent variables, their guarantees typically hold only up to equivalence classes and rely on restrictive structural and cardinality assumptions that are incompatible with Hawkes dynamics. For time-series data, LPCMCI (Gerhardus and Runge, 2020) adapts conditional-independence testing to temporal domains. However, it assumes weak autocorrelation and exogenous latent variables, assumptions that are violated in Hawkes processes where dense cross-lag dependencies and endogenous latent subprocesses naturally arise. Although our method also leverages second-order (rank) statistics, it differs in two key respects. First, it targets subprocesses in multivariate Hawkes systems rather than static variables and accommodates both endogenous and exogenous latent subprocesses. Second, it establishes identifiability through time-aware rank constraints specifically tailored to Hawkes dynamics, avoiding the infeasible assumptions underlying existing rank- or independence-test based approaches.
@@ 3 Partially Observed Multivariate Hawkes Process-based Causal Model
This section introduces the causal modeling framework for partially observed Hawkes processes and establishes the key definitions that support the subsequent results on structure discovery.
@@ 3.1 Multivariate Hawkes Process
A multivariate Hawkes process is a self-exciting point process modeling temporal dependencies among events via a set of counting subprocesses $\mathcal{N}_{\mathcal{G}}=\{N_{i}\}_{i=1}^{l}$, where $N_{i}(t)$ records the number of type-$i$ events up to time $t$ (Hawkes, 1971; Laub et al., 2015). For each $i\in\{1,\dots,l\}$, the intensity of subprocess $N_{i}$ that governs the event-triggering behavior is:
@@
where $\mu_{i}$ is the background intensity, and $\phi_{ij}(s)\geq 0,\forall s\in(0,\infty)$ is excitation function, which measures the decaying influence of historical type-$j$ events on the subsequent type-$i$ events and is piecewise continuous. (Strictly) stationarity requires the spectral radius of the influence matrix $\Phi\in\mathbb{R}^{l\times l}$ with entries $\Phi_{ij}=\int_{0}^{\infty}\phi_{ij}(s)ds$ to be less than one (Bacry and Muzy, 2016). A detailed exposition of Hawkes process is deferred to Appendix˜B. See also Fig. 1(a) for illustration.
We are interested in identifying, for each subprocess $N_{i}$, the minimal set of subprocesses $\mathcal{P}_{\mathcal{G}}\subseteq\mathcal{N}_{\mathcal{G}}$ such that $\lambda_{i}(t)$ depends only on the historical events of the subprocesses in $\mathcal{P}_{\mathcal{G}}$ and not on others. Formally, this corresponds to $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$ for each $N_{j}\in\mathcal{P}_{\mathcal{G}}$, and zero otherwise. In this case, $N_{i}$ is said to be locally independent (Didelez, 2008) of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$.
@@ 3.2 Model Definition
To formalize our framework, we define a graphical causal model for multivariate Hawkes processes, where nodes represent subprocesses and directed edges correspond to nonzero excitation functions. The goal is to recover the causal structure among both observed and latent subprocesses.
@@
Let $\mathcal{G}\coloneq(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}})$ be a directed graph, where each node $N_{i}\in\mathcal{N}_{\mathcal{G}}$ represents a subprocess in a multivariate Hawkes process. A directed edge $E_{ij}\in\mathcal{E}_{\mathcal{G}}$ exists iff $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$. The node set $\mathcal{N}_{\mathcal{G}}$ consists of $p$ observed nodes $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$ and $q$ latent nodes $\mathcal{L}_{\mathcal{G}}\coloneqq\{L_{i}\}_{i=1}^{q}$, which correspond to the observed and latent subprocesses, respectively.
The PO-MHP model naturally allows cycles and self-loops that are typically challenging to analyze (Claassen and Mooij, 2023). Two subprocesses $N_{i}$ and $N_{j}$ form a directed cycle if there exist directed paths from $N_{i}$ to $N_{j}$ and from $N_{j}$ back to $N_{i}$. Any subprocess $N_{i}$ has self loop if $\int_{0}^{t}\phi_{ii}(t-s)\,dN_{i}(s)>0$. Furthermore, it allows for the presence of latent subprocesses, and directed edges may exist between any pair of subprocesses, whether observed or latent; both $N_{i}$ and $N_{j}$ may be either observed or latent. To the best of our knowledge, it is the first principled work to investigate such a general structure in Hawkes processes.
Definition 3.2 (Cause and Effect).
For any $N_{i},N_{j}\in\mathcal{N}_{\mathcal{G}}$, if there exists a directed path from $N_{i}$ to $N_{j}$, then $N_{j}$ is said to be a effect of $N_{i}$, and $N_{i}$ is said to be a cause of $N_{j}$.
Definition 3.3 (Parent-Cause Set).
@@
Proposition 3.4 (Parent-Cause Set and Local Independence).
Subprocess $N_{i}$ is locally independent (defined in Section˜3.1) of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$ if and only if $\mathcal{P}_{\mathcal{G}}$ is the parent-cause set of $N_{i}$ in $\mathcal{N}_{\mathcal{G}}$.
Proposition˜3.4 is the PO-MHP analogue of the local Markov property: even with cycles and self-loops, conditioning on the parent-cause set $\mathcal{P}_{\mathcal{G}}$ screens off $N_{i}$ from all other subprocesses.
Figure 1: Figure 1: Illustration of multivariate Hawkes processes. (a) Point process representation with three subprocesses $N_{1},N_{2},N_{3}$, where the continuous timeline is partitioned into intervals of length $\Delta$. (b) The corresponding summary causal graph, the central object of this paper, with causal relations $N_{1}\leftarrow N_{2}\leftrightarrow N_{3}$ and self-loops on all nodes. (c) The window causal graph, showing the underlying time-lagged causal mechanism: each node denotes the count in one interval of length $\Delta$, modeled as a weighted sum of lagged parent nodes plus noise (Eq. 1). (d) A minimal example with a latent subprocess $L_{1}$ confounding $O_{1}$ and $O_{2}$, highlighting the primary focus of this paper.
@@ 4 Structure Identification in Partially Observed Hawkes Processes
In this section, we formalize partially observed Hawkes processes in a discrete-time framework and develop rank-based tests on observed counts that enable identification of the summary causal graph, including relationships mediated by latent subprocesses.
@@ 4.1 From Continuous-Time to Discrete-Time Representation
Directly inferring the causal structure among subprocesses in a multivariate Hawkes process is challenging, particularly with latent subprocesses, as Eq. 1 defines a continuous-time stochastic process where $\lambda_{i}(t)$ is the expected history-dependent instantaneous rate. Instead of relying on conventional maximum-likelihood fitting, we present an explicit correspondence between Hawkes dynamics and a specified discrete-time linear autoregressive causal model, which enable the identification of latent subprocesses and the causal structure by applying statistical tests only on observed count data.
Theorem 4.1 (Hawkes Process as a Linear Autoregressive Model).
Let $\mathcal{N}_{\mathcal{G}}\coloneq\{N_{i}\}^{l}_{i=1}$ be a stationary multivariate Hawkes process with background intensities $\{\mu_{i}\}_{i=1}^{l}$ and excitation functions $\{\phi_{ij}(s)\}_{i,j=1}^{l}$. Define the discretized event count in the $n$-th time window of size $\Delta\in(0,\delta)$ as
@@
where $\delta>0$ depends on the moment structure of the process. Then, as $\Delta\rightarrow 0$, the Hawkes process admits the linear autoregressive representation
@@
where $\theta^{(0)}_{i}=\Delta\cdot\mu_{i}$ is the background parameter, $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)ds$ is the excitation coefficient, and $\varepsilon_{i}^{(n)}$ denotes the $n$-th realization of a serially uncorrelated white noise sequence.
Theorem˜4.1 shows that each current-bin count $N_{i}^{(n)}$ (referred to as variable hereafter) is a linear combination of lagged counts $\{N_{j}^{(n-k)}\}_{j\in\{1,\dots,l\}}^{k\in\{1,\dots,n\}}$ plus noise. The discretized variables therefore encode the causal structure of the underlying continuous-time subprocesses. As illustrated in Fig. 1, a directed edge $N_{2}\!\rightarrow\!N_{1}$ in the summary graph corresponds to edges from all lagged variables of $N_{2}$ into $N_{1}^{(n)}$ in the window graph. This summary–window correspondence allows us to infer the causal structure among subprocesses by testing structural relations among the discretized variables. The proof of Theorem˜4.1 appears in Appendix˜G.
In practice, it is unnecessary to include all lags $k=1,\dots,n$. Because the excitation function $\phi_{ij}(s)$, which serves as a decay kernel, typically has finite support, the coefficients $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)\,ds$ vanish for large $k$, distant lags $N_{j}^{(n-k)}$ therefore have little influence on $N_{i}^{(n)}$. We define the smallest cutoff $K$ at which this occurs as the number of effective lags. Accordingly, we truncate to at most $m$ lags with $m\!\geq\!K$, a standard finite-window practice in time-series estimation (Hyvärinen et al., 2010; Peters et al., 2013), though they assume acyclic summary graphs and understate latent components. A practical way to estimate $K$ is to retain lagged variables whose correlations with current variables remain statistically significant. Moreover, a bin width $\Delta$ small relative to the support of the excitation function is sufficient; see Section˜Q.3 for an empirical illustration.